OpenStudy (anonymous):

Sara is twice as tall as Willie ans 2/3 as tall as Andre.If the average height of all three is 4 ft.how tall is Andre? a.2ft b.4ft c.5ft d.ft e.8ft

OpenStudy (anonymous):

Well, make Sara's height S, Willie's height W, and Andre's height A. I'm assuming that you meant Sara is twice Willie's height AND 2/3 Andre's height. This means: S= 2*W and S=2/3*A This means that W=S/2=2/3*A/2=A/3 and that S=2/3*A The average of their heights is thus (S+W+A)/3=(A/3+2/3*A+A)/3 or 2A/3 Since the average height is 4 ft, 2A/3=4 feet. Solve for A: A=6 feet. I don't see that as an option, but I'm guessing d. was 6 feet, which would be your answer.

OpenStudy (anonymous):

4ft

OpenStudy (anonymous):

S=2W S=(2/3)A HENCE W = S/2 A = (3/2)S AVERAGE OF ALL IS 4 HENCE (W+A+S)/3 = 4 ((S/2)+((3/2)S)+S)/3=4 (S+3S+2S)/6=4 6S/6=4 So S=4 W=8 A =6

OpenStudy (anonymous):

/\ rahulpn25, If W=S/2, then how does W=8 when S=4? Using your method, S=4, so W=2,. Whatever the case, you still come out with A=6, which is what they're looking for. They're looking for ANDRE's height, not Sara's. Therefore, as you proved, andre's height is 6 feet.